Construct the indicated confidence interval for the population mean mu using the t-distribution. Assume the population is normally distributed. c = 0.99, x̅ =12.1, s = 4.0, n = 7
Answer #1

Solution :

t/2,df = 3.707

Margin of error = E = t/2,df* (s /√n)

= 3.707 * (4 / √7)

Margin of error = E = 5.6

The 99% confidence interval estimate of the population mean is,

- E + E

12.1 - 5.6

6.5

(6.5 , 17.7)